ELECTRONIC SPECTRA OF THE JET-COOLED ACETAMINOPHEN

\underline{SEUNG JUN LEE}, AHREUM MIN, YUSIC KIM and MYONG YONG CHOI, Department of Chemistry (BK21) and Research Institute of Natural Science, Gyeongsang National University, Jinju 660-701, Korea; JINYOUNG CHANG, SANG HAK LEE, Department of Chemistry, Seoul National University, Seoul 151-747, Korea; and SEONG KEUN KIM, Department of Chemistry,and Department of Biophysics and Chemical Biology, Seoul National University, Seoul 151-747, Korea.

Resonant two-photon ionization (R2PI), laser induced fluorescence (LIF) and UV-UV double resonance spectra of the jet-cooled acetaminophen, widely used as a pain reliever and fever reducer, were obtained in the gas phase. Conformational characterizations for acetaminophen will be presented with an aid of spectroscopic techniques and DFT B3LYP calculations.